1) Given the following cost information for a price-taking firm, what is the maximum amount of...

1)

Given the following cost information for a price-taking firm, what is the maximum amount of profit that can be earned if the price is $60?

Quantity 0 1 2 3 4 5 6
Total Cost 50 100 130 150 190 250 330

      

Group of answer choices

a $300

b $40

c $30

d $50

2)

Given the following cost information for a price-taking firm, what is the profit-maximizing quantity given a market price of $40?

Quantity 0 1 2 3 4 5 6 7 8
AVC 100 75 60 50 48 50 54.3 62.5

         

Group of answer choices

None of these answers

a Q = 6

b Q = 5

c Q = 8

d Q = 7
Answer #1

1) Profit Maximization happens when MC=MR and since Price is constant for the outputs MR=P=$60

Fixed cost=$50 i.e. the cost at no output

MC for unit 1=100-50=$50,unit 2=130-100=$30, unit3=150-130=$20, unit4=190-150=$40,unit5=250-190=$60,unit6=330-250=$80.

Since MC=MR is at unit5 which is $60, Profit Maximization will happen till unit 5 and Profit=Total Revenue-Total cost=60*5-250=$50 option d) is correct

2) Total variable cost= AVC*Quantity Hence, for output1=100*1=$100,MC=$100, Output 2=75*2=$150,MC=$150-100=$50, output 3=60*3=$180,MC=$180-$150=$30, output 4==50*4=$200,MC=200-180=$20, output 5=48*5=$240,MC=240-200=$40, output 6=50*6=300,MC=300-240=$60, output 7=54.3*7=380.1,MC=380.1-300=$80.1, output 8=62.5*8=$500,MC=500-380.1=119.9

Since MC=P=$40 is at output 5 Hence, profit maximizing output is Q=5 option b